OpenStudy (anonymous):

I need to find an equation of the line containing the given pair of points (3,4) and (9,9). I have figured out that the slope is 5/6 but I can not figure out the b of y=5/6+b.

OpenStudy (anonymous):

First of all, it's 5/6x +b Second, use point-slope form to get the answer : y - y1 = m(x-x1) y-9 = 5/6*(x-9) y = 5/6x + 1.5

OpenStudy (anonymous):

instead try this: given points: (x1,y1) and (x2,y2) (x-x1)/(x2-x1)=(y-y1)/(y2-y1) substituting, (x-3)/(9-3)=(y-4)/(9-4) (x-3)/6=(y-4)/5 5x-15=6y-24 5x-6y+9 is the equation of the line u can check it. the slope calculated from this equation is same as the one u calculated